# Exploring the Influence of Information Entropy Change in Learning Systems

### v1

@@ 4.5 Domain Adaption Results
Unsupervised domain adaptation (UDA) aims to learn transferable knowledge across the source and target domains with different distributions [41] [62]. Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home [61] and VisDA2017 [42]. Detailed datasets introduction and experiments training settings are in the supplementary. The objective function is borrowed from TVT [66], which is the first work that adopts Transformer-based architecture for UDA. The results are shown in Table 6 and 7. The ViT-B with positive noise achieves better performance than the existing works. These results show that positive noise can improve model generality, therefore, benefit deep models in domain adaptation tasks.
@@ A.1.1 Add Gaussian Noise to Original Images
$\begin{split}MI(\mathcal{T},\boldsymbol{\epsilon})=&H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X+\epsilon})\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X+\epsilon}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\end{split}$ (43)
@@
$\left\{\begin{matrix}MI(\mathcal{T},\boldsymbol{\epsilon})>0&if\ \frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}>1\\ MI(\mathcal{T},\boldsymbol{\epsilon})\leq 0&if\ \frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\leq 1\end{matrix}\right.$ (44)
@@
As a result, $\frac{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}{|\boldsymbol{\Sigma_{Y|X}}|}$ is expressed as:
@@
Combine Equations 57 and 2 together, the mutual information is expressed as:
$\begin{split}MI(\mathcal{T},\boldsymbol{\epsilon})=&\frac{1}{2}\log\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{X_{i}}}\mathrm{cov}^{2}(X_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))})}\\ =&\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}\end{split}$ (58)
@@
To determine if Gaussian noise can be positive noise, we need to determine whether the mutual information is large or smaller than 0:
@@
Combine the Equations 44 and 61, we can get the conclusion:
$\left\{\begin{matrix}MI(\mathcal{T},\boldsymbol{\epsilon})>0&\mathrm{if}\ f(\sigma_{\epsilon}^{2})>0\\ MI(\mathcal{T},\boldsymbol{\epsilon})\leq 0&\mathrm{if}\ f(\sigma_{\epsilon}^{2})\leq 0\end{matrix}\right.$ (62)
From the above equations, the sign of the mutual information is determined by the statistical properties of the data samples and labels. Since ${\epsilon}^{2}\geq 0$ and $\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}\geq 0$, we have a deep dive into the residual part, i.e.,
@@
where $\rho_{X_{i}Y_{i}}$ is the correlation coefficient, and $\rho^{2}_{X_{i}Y_{i}}\in[0,1]$. As a result, the sign of the mutual information in the Gaussian noise case is negative. We can conclude that Gaussian noise added to the images is harmful to the task.
@@ A.1.2 Inject Gaussian Noise in Latent Space
Thus, the mutual information of injecting Gaussian noise can be formulated as:
$\begin{split}MI(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})\end{split}$ (65)
@@
$\begin{split}MI(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{X}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{X+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\\ =&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}})}\end{split}$ (66)
@@ A.2.1 Add Linear Transform Noise to Original Images
$\begin{split}MI(\mathcal{T},Q\boldsymbol{X})=&H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X}+Q\boldsymbol{X})\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X}}|-\frac{1}{2}\log|{\Sigma_{\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X}}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|{\Sigma_{\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X}}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}}\boldsymbol{\Sigma^{-1}_{X}}\boldsymbol{\Sigma_{XY}}|}\\ =&0\end{split}$ (71)
The mutual information of 0 indicates that the implementation of linear transformation to the original images could not reduce the complexity of the task.
@@ A.2.2 Inject Linear Transform Noise in Latent Space
The mutual information of injecting linear transform noise can be formulated as:
$\begin{split}MI(\mathcal{T},Q\boldsymbol{X})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}+Q\boldsymbol{X})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+Q\boldsymbol{X})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma}_{\boldsymbol{X}}||\boldsymbol{\Sigma_{Y}}-\boldsymbol{\Sigma_{YX}}{\boldsymbol{\Sigma_{X}^{-1}}}\boldsymbol{\Sigma_{XY}}|}{|\boldsymbol{\Sigma}_{(I+Q)\boldsymbol{X}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}}\boldsymbol{\Sigma^{-1}_{X}}\boldsymbol{\Sigma_{XY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (72)
Since we want the mutual information to be greater than 0, we can formulate Equation 72 as an optimization problem:
$\begin{split}&\max_{Q}MI(\mathcal{T},Q\boldsymbol{X})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (73)
@@
    where $\beta$ can be at a random position beside the diagonal. As we can see from the simple example, Row Sum breaks the fourth constraint and make mutual information smaller than 0.
@@
However, if we combine the elementary row operations, it is possible to make the mutual information greater than 0 as well as satisfy the constraints. For example, we combine the Row Swap and Scalar Multiplication to generate the $Q$:
@@
In this case, $MI(\mathcal{T},Q\boldsymbol{X})>0$ when $Q=-0.5I$. The constraints are satisfied. This is just a simple case that adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise. To this end, we demonstrate that linear transform can be positive noise.
@@ A.3.1 Add Salt-and-pepper Noise to Original Images
Therefore, the mutual information with salt-and-pepper noise is:
$MI(\mathcal{T},\boldsymbol{\epsilon})=H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X}\boldsymbol{\epsilon})=0$ (83)
@@ A.3.2 Add Salt-and-pepper Noise in Latent Space
The mutual information of injecting salt-and-pepper noise can be formulated as:
$\begin{split}MI(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})\\ =&-\sum_{\boldsymbol{X}\in\mathcal{X}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{X},\boldsymbol{Y})\log p(\boldsymbol{X},\boldsymbol{Y})-\sum_{\boldsymbol{X}\in\mathcal{X}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}\sum_{\boldsymbol{\epsilon}\in\mathcal{E}}p(\boldsymbol{X\epsilon},\boldsymbol{Y})\log p(\boldsymbol{X\epsilon},\boldsymbol{Y})\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X\epsilon},\boldsymbol{Y})}\right]\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-H(\epsilon)\end{split}$ (84)
The mutual information is smaller than 0, therefore, the salt-and-pepper is a pure detrimental noise to the latent representations.

### v2

@@ 4.5 Domain Adaption Results
Unsupervised domain adaptation (UDA) aims to learn transferable knowledge across the source and target domains with different distributions [41] [62]. Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home [61] and VisDA2017 [42]. Detailed datasets introduction and experiments training settings are in the supplementary. The objective function is borrowed from TVT [66], which is the first work that adopts Transformer-based architecture for UDA. The results are shown in Table 6 and 7. The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results demonstrate that positive noise also works in domain adaptation tasks, where out-of-distribution (OOD) data exists.
@@ A.1.1 Add Gaussian Noise to Original Images
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X+\epsilon})\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X+\epsilon}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\end{split}$ (42)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&if\ \frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}>1\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&if\ \frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\leq 1\end{matrix}\right.$ (43)
@@
As a result, $\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}$ is expressed as:
@@
therefore, the entropy change is expressed as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&\frac{1}{2}\log\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{X_{i}}}\mathrm{cov}^{2}(X_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))})}\\ =&\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}\end{split}$ (57)
@@
To determine if Gaussian noise can be positive noise, we need to determine whether the entropy change is large or smaller than 0:
@@
Combine the Equations 43 and 60, we can get the conclusion:
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&\mathrm{if}\ f(\sigma_{\epsilon}^{2})>0\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&\mathrm{if}\ f(\sigma_{\epsilon}^{2})\leq 0\end{matrix}\right.$ (61)
From the above equations, the sign of the entropy change is determined by the statistical properties of the data samples and labels. Since ${\epsilon}^{2}\geq 0$ and $\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}\geq 0$, we have a deep dive into the residual part, i.e.,
@@
where $\rho_{X_{i}Y_{i}}$ is the correlation coefficient, and $\rho^{2}_{X_{i}Y_{i}}\in[0,1]$. As a result, the sign of the entropy change in the Gaussian noise case is negative. We can conclude that Gaussian noise added to the images is harmful to the task.
@@ A.1.2 Inject Gaussian Noise in Latent Space
Thus, the entropy change of injecting Gaussian noise can be formulated as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})\end{split}$ (64)
@@
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{X}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{X+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\\ =&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}})}\end{split}$ (65)
@@ A.2.1 Add Linear Transform Noise to Original Images
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})=&H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X}+Q\boldsymbol{X})\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X}}|-\frac{1}{2}\log|{\Sigma_{\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X}}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|{\Sigma_{\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X}}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}}\boldsymbol{\Sigma^{-1}_{X}}\boldsymbol{\Sigma_{XY}}|}\\ =&0\end{split}$ (70)
The entropy change of 0 indicates that the implementation of linear transformation to the original images could not reduce the complexity of the task.
@@ A.2.2 Inject Linear Transform Noise in Latent Space
The entropy change of injecting linear transform noise can be formulated as:
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}+Q\boldsymbol{X})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+Q\boldsymbol{X})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma}_{\boldsymbol{X}}||\boldsymbol{\Sigma_{Y}}-\boldsymbol{\Sigma_{YX}}{\boldsymbol{\Sigma_{X}^{-1}}}\boldsymbol{\Sigma_{XY}}|}{|\boldsymbol{\Sigma}_{(I+Q)\boldsymbol{X}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}}\boldsymbol{\Sigma^{-1}_{X}}\boldsymbol{\Sigma_{XY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (71)
Since we want the entropy change to be greater than 0, we can formulate Equation 71 as an optimization problem:
$\begin{split}&\max_{Q}\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (72)
@@
    where $\beta$ can be at a random position beside the diagonal. As we can see from the simple example, Row Sum breaks the fourth constraint and make entropy change smaller than 0.
@@
However, if we combine the elementary row operations, it is possible to make the entropy change greater than 0 as well as satisfy the constraints. For example, we combine the Row Swap and Scalar Multiplication to generate the $Q$:
@@
In this case, $\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})>0$ when $Q=-0.5I$. The constraints are satisfied. This is just a simple case that adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise. To this end, we demonstrate that linear transform can be positive noise.
@@ A.3.1 Add Salt-and-pepper Noise to Original Images
Therefore, the entropy change with salt-and-pepper noise is:
$\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X}\boldsymbol{\epsilon})=0$ (82)
@@ A.3.2 Add Salt-and-pepper Noise in Latent Space
The entropy change of injecting salt-and-pepper noise can be formulated as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})\\ =&-\sum_{\boldsymbol{X}\in\mathcal{X}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{X},\boldsymbol{Y})\log p(\boldsymbol{X},\boldsymbol{Y})-\sum_{\boldsymbol{X}\in\mathcal{X}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}\sum_{\boldsymbol{\epsilon}\in\mathcal{E}}p(\boldsymbol{X\epsilon},\boldsymbol{Y})\log p(\boldsymbol{X\epsilon},\boldsymbol{Y})\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X\epsilon},\boldsymbol{Y})}\right]\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-H(\epsilon)\end{split}$ (83)
The entropy change is smaller than 0, therefore, the salt-and-pepper is a pure detrimental noise to the latent representations.